OpenStudy (anonymous):

What is the slope of a line that is perpendicular to the line whose equation is 2y = 3x - 1? a) -3 b) -1/3 c) -2/3

OpenStudy (godlovesme):

first isolate the y

OpenStudy (anonymous):

What is your understanding of this question so far? We can work from there :)

OpenStudy (godlovesme):

@makayla.boze u there? o_O

OpenStudy (pawanyadav):

-2/3 as product of slope of two perpendicular line is -1

OpenStudy (pawanyadav):

So 3/2×m=-1 m=-1÷3/2=-2/3